# Teaching Cloud Infrastructure and Scalable Application Deployment in an Undergraduate Computer Science Program

### v1

@@ Abstract.
Making successful use of cloud computing for deploying scalable web applications requires nuanced approaches to both system design and deployment methodology, involving reasoning about the elasticity, cost, and security models of cloud services. Students commonly interact with cloud abstractions early in their technical careers, including during internships and academic research. Building cloud-native applications without a firm understanding of the fundamentals of cloud engineering can leave students susceptible to cost and security pitfalls common to cloud platforms. Yet, cloud computing is not commonly taught at the undergraduate level, because the technology and practices behind modern cloud deployment, such as containerization and infrastructure-as-code (IaC), have only recently matured into a set of general principles independent from specific providers’ offerings. To address this gap, we designed an undergraduate-level course around these principles that framed cloud infrastructure deployment as a software engineering practice in support of scalable web applications, emphasizing the value of both cloud deployment and application design skills in building robust cloud-native systems. Our course featured a number of hands-on assignments that gave students experience with modern, best-practice concepts and tools such as IaC, containerization, observability, serverless computing, and continuous integration and deployment. We describe the design of the course, our experience teaching its initial offering in Winter 2024, and provide our reflections on what worked well and potential areas for improvement. Our course material is publicly available at https://infracourse.cloud.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the 56th ACM Technical Symposium on Computer Science Education; February 26–March 1, 2025; Pittsburgh, PAccs: Social and professional topics Computing educationccs: Software and its engineering Cloud computing
@@ 1. Introduction
The use of cloud computing and infrastructure to deploy user-facing applications and to perform computing-intensive tasks is more important today than ever. Organizations large and small in industry and academia alike are increasingly relying on cloud infrastructure for their computing workloads in order to better scale applications or run high-performance experiments, such as machine learning training and inference, with less engineering and maintenance overhead (Goodwins, 2015).
@@
Making optimal use of cloud infrastructure is a nontrivial engineering problem. Cloud-native development and deployment diverges from traditional, on-premise practices in a number of key ways that together represent a paradigm shift in the mindset required to approach such tasks. For example, modern cloud-deployed applications often run on compute platforms where state is ephemeral, such as managed container clusters, and thus must be designed to store state in separate managed services such as databases and object storage (McCance, 2012). Cloud infrastructure also gives developers tools to more finely and elastically control the compute and storage capacity needed to support their applications, saving cost when resources are not in use while improving scalability during demand spikes (Lehrig et al., 2015). Securing cloud-deployed applications also requires understanding a different set of abstractions. Though such workloads run on third-party hardware, cloud providers only assume security responsibility for the cloud platform and its underlying hardware and software infrastructure, while delegating the responsibility of how cloud services are used to developers (Services, 2024; Cloud, 2023; Azure, 2023).
@@
Working with cloud platforms is tricky and requires a nuanced understanding of these concepts. Many engineers often experience difficulties stemming from a poor understanding of security, elasticity, and maintenance, with potentially severe costs. For example, in late 2023, attackers exploited leaked AWS API keys to mine cryptocurrency within victims’ AWS accounts, leading to thousands of dollars in unexpected costs (McCloskey and Tafani-Dereeper, 2024). Another issue can stem from choosing the wrong data storage service for a given application, which can lead to poor scalability and high data transfer expenses.
@@
Students often encounter a need to heavily interact with cloud abstractions early in their careers, both within the university context and in industry. Many CS curricula in universities including our institution feature capstone projects where students work cooperatively on a complex web application (Li et al., 2023; Shakil and Denny, 2024), where deployment is often a required component. Instruction on doing so is typically limited and relegated to the end of the course, mostly focusing on deploying to a single virtual machine (VM) rather than meaningfully interacting with modern, managed cloud abstractions at both the deployment and application level. Similarly, during software engineering internships, students often work with existing cloud-deployed services such as containers in clusters or even deploy new services themselves.
@@
Effectively preparing students for these roles requires a more comprehensive primer on working with cloud services in order to help them avoid common pitfalls. Although there are a number of existing university courses that cover cloud computing, most are offered at a graduate or professional level and focus largely on the engineering and research behind cloud platforms themselves (Asanovic and Karandikar, 2023; Kozyrakis, 2024; Birman, 2022). Cloud deployment is not commonly taught at the undergraduate level, as cloud infrastructure offerings went through a period of rapid change in the past and often diverged drastically between providers (Kroonenburg, [n. d.]). But in recent years, there has been a maturation of the technologies and practices involved in deployment beyond provider-specific tools. These concepts, such as infrastructure-as-code (IaC), containerization and container orchestration, functions-as-a-service (FaaS), and other managed abstractions on top of open-source compute and data services, are also now more relevant within a software engineering career as the separation between engineering and operations has narrowed (Buchanan, 2024).
Accordingly, we developed a new undergraduate course at our institution known as ”Cloud Infrastructure and Scalable Application Deployment.” Our course focused on building student fluency with these cloud infrastructure-as-a-service (IaaS) concepts and tools in a hands-on manner, treating cloud deployment as a software engineering practice.
@@
We achieved this goal through a series of lectures that placed significant emphasis on application design in a way that best takes advantage of modern cloud services for scalability to a broad number of users with minimal cost and system administration overhead. Our course additionally offered students a safe sandbox in which to experiment with cloud deployments, while providing some structure in learning how to do so systematically. This involved designing a novel set of assignments that gave students practice in deploying applications in modern patterns drawing on complex managed services while using IaC.
@@
We present insights from our experience teaching the inaugural offering of the course in Winter 2024. We hope that our reflections and course materials can be useful to educators at other institutions interested in teaching students practical cloud computing skills.
@@ 2. Related Work
Usage of the cloud also features in a number of other CS courses, even if it is not the main focus. Many universities offer courses on DevOps, in which students learn about the tools and practices that enable rapid software development and deployment (Benni et al., 2019; Alves and Rocha, 2021; Fernandes et al., 2022; Verdicchio, 2023). Cloud infrastructure is a key part of DevOps, as it enables the automation of pipelines for building, testing, and deploying software. However, these courses mostly focus on the processes that make such a software lifecycle efficient, rather than the infrastructure that supports both these processes and the deployed applications.
@@ 3. Course Overview and Logistics
We designed a course entitled ”Cloud Infrastructure and Scalable Application Deployment” and offered it as an undergraduate elective at our U.S. institution in Winter 2024. Our institution follows a quarter system, so the course was taught over 10 weeks.
@@ 3.1. Learning Objectives
    Understand how to design cloud-native applications to take advantage of the elasticity, cost, and system administration benefits provided by the building blocks of cloud infrastructure.
@@
3.2. Class Makeup
In total, 50 students were enrolled in the course until the end of the term, including 15 women and 7 students from underrepresented minority backgrounds. Most students were either juniors, seniors, or graduate students, although there were a limited number of underclassmen enrolled. Although the course was open to students of all majors (by institutional policy), most students were CS majors.
We intentionally only required programming maturity up to an introductory systems class and some experience using command-line tools as a prerequisite in order to keep the course accessible to underclassmen. However, of the 34 students that completed an intro survey and remained enrolled in the class until the end, 76% had completed an operating systems course and 41% had completed a web or mobile development course. Additionally, 79% had some experience with backend or frontend web development, 68% had experience using an IaaS cloud provider, and 35% had experience with containerization. The first offering of the course was a pilot that more experienced students likely self-selected into, and we expect that future offerings will have a more diverse set of students with less prior experience.
@@ 3.3. Course Structure
Over 10 weeks, the course met for 80 minutes twice a week. Our lectures covered the core areas of cloud application deployment, detailing the types of services offered by public cloud providers and the considerations behind selecting and deploying these services for a given application. We additionally spent significant time on discussing system and application design for scalability, security, and cost efficiency.
There were four hands-on programming assignments over the course of the quarter, each worth 15% of a student’s final grade, which we describe in Section 4.2. Assignments focused on systematically deploying a pre-built application to the cloud in modern best-practice patterns using IaC. We encouraged students to work in pairs on these assignments, and students were allowed to discuss conceptual strategies with other students in the class as long as they refrained from sharing code.
Lastly, the course featured a final project, worth 40% of a student’s final grade, where students were given the opportunity to deploy an application of their own choosing to the cloud.
@@ 4. Course Content
Our course aimed to present a set of fundamental practices necessary for publishing and managing usable web applications on the public cloud as well as application design that takes advantage of cloud abstractions. Our material, including lectures, assignments, and the infrastructure we used to manage the class, is open-source and publicly available online.
@@ 4.1. Lecture Content
The first half of our course discussed the building blocks of cloud infrastructure that help in deploying web applications. Topics covered during this phase of the course included the basics of compute, network, and storage resources on the cloud, containers and container orchestration, and serverless and edge computing using Functions-as-a-Service (FaaS).
@@
Security was a significant focus area within the course. We emphasized the shared responsibility model of cloud security, where providers take responsibility for securing the hardware and software underlying cloud platforms whereas developers take responsibility for how they use cloud services. We included a dedicated lecture on identity and access management (IAM) and its high attack surface in the context of modern serverless and containerized workloads running on the cloud. For all other lectures, where relevant, we made sure to include security considerations and best practices to minimize the attack surface of the services we discussed.
We covered all conceptual material and general concepts in lectures in a cloud provider-agnostic function, but demonstrated specific examples on Amazon Web Services (AWS) where required. We summarize the content covered in these lectures in Table 1.
@@ 4.1.1. Guest Lectures
We also invited four guest lecturers from industry to provide additional insight and depth on select topics. All speakers were deeply involved with leading the design and deployment of cloud-native applications at their respective employers. The content they covered is summarized in Table 1.
@@ 4 Continuous Integration & Deployment, Container Registries
Our course featured four hands-on technical assignments that were designed to give students experience with the tools and services discussed in lectures. Our assignments focused on deploying web services to AWS, which we chose due to its predominant market share and the maturity of its cloud service offerings. AWS additionally granted each student \$200 in credits for the duration of the course, which covered students’ cloud resource costs. However, we emphasize that these assignments can be generalized to any cloud provider and are not AWS-specific.
@@ 4.3. Final Project
The final project for the course was an open-ended assignment in which students were tasked with deploying a web application of their choosing to the public cloud. Students were allowed to choose either an application they designed themselves or an existing open-source project. The project was a way for students to demonstrate their ability to cumulatively apply the concepts they learned without the scaffolding we provided in the assignments. In particular, it gave students the opportunity to architect a new deployment rather than merely instantiating cloud resources for an existing architecture.
As such, students were expected to be able to tie together multiple cloud services using IaC to create an elastic deployment that could scale with demand and followed security best practices. Deliverables included the source code for the IaC used to deploy the application, a brief write-up of the deployment’s architecture and services used, and a link to the deployed application for evaluation.
@@ 5. Discussion
In total, 39 students filled out the survey. 72% of students stated they learned ”a great deal” or ”a lot” (a score of 4 or 5 on a 1-5 scale). Students broadly found value in the practical skills they learned: a few students reported that the course built up ”actual confidence in my ability to use the cloud intelligently,” and that ”this was the missing building block in building full-stack cloud awareness.” Students also overwhelmingly reported that they found the course useful for their future careers.
We also received feedback that the course was harder and more time-consuming than students expected. We believe this was partially due to the sheer number of new concepts, services, and tools that students were exposed to, as well as general teething challenges common to any first offering of a course. However, many issues also stemmed from our miscalibration of assignment difficulty with the time we gave students to complete them, the scaffolding we provided in assignment starter code that prevented students from making more architectural choices, and state management issues with the AWS CDK that led to significant challenges in debugging deployments. We are currently in the process of addressing these issues based on student feedback for the next offering of the course.
@@ 5.1.1. Course infrastructure
In order to easily manage a course of 50 students with only three staff, we built a significant amount of custom infrastructure to support students given the variety of resources they interacted with throughout the course.
@@ 5.2.1. Building domain-specific debugging skills
During our pilot offering, we erred on the side of providing students with more scaffolding for IaC deployments. Unfortunately, this made it more difficult for students to practically apply application design concepts they learned in lecture. In the future, we intend to give students more architectural freedom for their submissions. We envision grading some assignments based on performance and reliability rather than on the IaC itself, which may help in making assignments more open-ended without much additional overhead.
@@ 6. Conclusion
We present our experience teaching an undergraduate course on cloud infrastructure and scalable application deployment. Students learned practical skills related to deploying and managing cloud infrastructure, as well as application design that takes advantage of managed cloud services. Our course featured a number of hands-on assignments that gave students experience with modern, best-practice concepts such as IaC, containerization, and serverless computing. While there are several improvements we can make to future offerings, both students and course staff felt that the course was successful in building student fluency with cloud fundamentals and tools useful for launching scalable applications and for further career growth. We hope that our experiences and public course materials can be useful to educators interested in addressing the gap in practical cloud computing education for undergraduate students.

### v2

@@ Abstract.
Making successful use of cloud computing requires nuanced approaches to both system design and deployment methodology, involving reasoning about the elasticity, cost, and security models of cloud services. Building cloud-native applications without a firm understanding of the fundamentals of cloud engineering can leave students susceptible to cost and security pitfalls. Yet, cloud computing is not commonly taught at the undergraduate level. To address this gap, we designed an undergraduate-level course that frames cloud infrastructure deployment as a software engineering practice. Our course featured a number of hands-on assignments that gave students experience with modern, best-practice concepts and tools including infrastructure-as-code (IaC). We describe the design of the course, our experience teaching its initial offering, and provide our reflections on what worked well and potential areas for improvement. Our course material is available at https://infracourse.cloud.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 56th ACM Technical Symposium on Computer Science Education V. 1; February 26-March 1, 2025; Pittsburgh, PA, USAbooktitle: Proceedings of the 56th ACM Technical Symposium on Computer Science Education V. 1 (SIGCSE TS 2025), February 26-March 1, 2025, Pittsburgh, PA, USAdoi: 10.1145/3641554.3701870isbn: 979-8-4007-0531-1/25/02ccs: Social and professional topics Computing educationccs: Software and its engineering Cloud computing
@@ 1. Introduction
Modern organizations are increasingly relying on cloud infrastructure for their computing workloads to better scale applications or run high-performance experiments.
@@
Making optimal use of cloud infrastructure is a nontrivial engineering problem. Cloud-native development and deployment diverges from traditional, on-premise practices in a number of key ways that together represent a paradigm shift in the mindset required to approach such tasks. For example, modern cloud-deployed applications often run on compute platforms where state is ephemeral and thus must be designed to store state in separate managed services (McCance, 2012). Cloud infrastructure gives developers tools to more finely control compute and storage capacity needed to support their applications (Lehrig et al., 2015). Securing cloud-deployed applications also requires understanding a different set of abstractions. Though such workloads run on third-party hardware, cloud providers only assume security responsibility for the cloud platform, while delegating the responsibility of how cloud services are used to developers (Services, 2024; Cloud, 2023; Azure, 2023).
@@
Effectively and securely working with cloud platforms is tricky and requires a nuanced understanding of cloud-native application design and deployment concepts. Many engineers experience difficulties stemming from a poor understanding of security, elasticity, and maintenance, with potentially severe costs. For example, in late 2023, attackers exploited leaked AWS API keys to mine cryptocurrency within victims’ AWS accounts, leading to thousands of dollars in unexpected costs (McCloskey and Tafani-Dereeper, 2024). Other issues can stem from choosing the wrong data storage service, which can lead to poor scalability and high data transfer expenses.
@@
Students often encounter cloud abstractions early in their careers, both within the university context and in industry. Many CS curricula in universities, including our institution, feature capstone projects where students work cooperatively on a complex web application (Li et al., 2023; Shakil and Denny, 2024), where deployment is often a required component. Instruction on doing so is typically limited and relegated to the end of the course, mostly focusing on deploying to a single virtual machine (VM) rather than meaningfully interacting with modern, managed cloud abstractions. Similarly, during software engineering internships, students often work with cloud-deployed services such as containers in clusters or deploy new services themselves.
@@
Effectively preparing students for these academic and industry roles requires a more comprehensive primer on working with cloud services. Although there are a number of existing courses that cover cloud computing, most are offered at a graduate or professional level and focus largely on the engineering and research behind cloud platforms themselves (Asanovic and Karandikar, 2023; Kozyrakis, 2024; Birman, 2022), rather than on deploying and designing cloud-native applications. Cloud deployment is not commonly taught at the undergraduate level, as cloud infrastructure offerings went through a period of rapid change in the past and often diverged drastically between providers (Kroonenburg, [n. d.]). But, in recent years, there has been a maturation of the technologies and practices involved in deployment beyond provider-specific tools. These concepts, such as infrastructure-as-code (IaC), containerization and container orchestration, functions-as-a-service (FaaS), and other managed abstractions on top of open-source compute and data services, are also now more relevant within a software engineering career as the separation between engineering and operations has narrowed (Buchanan, 2024).
@@
Accordingly, we developed a new undergraduate course known as ”Cloud Infrastructure and Scalable Application Deployment,” which focuses on building student fluency with these cloud infrastructure as a service (IaaS) concepts and tools in a hands-on manner. Our course consists of lectures that emphasize application design in a way that best takes advantage of modern cloud services. Our course additionally offered students a safe sandbox in which to experiment with cloud deployments, while providing structure in learning how to do so systematically. This involved designing a novel set of assignments that gave students practice in deploying applications in modern patterns using IaC.
@@
We present insights from our experience teaching the inaugural offering of the course in Winter 2024. We hope that our reflections and course materials can be useful to educators at other institutions.
@@ 2. Related Work
Usage of the cloud also features in a number of other CS courses, even if it is not the main focus. Many universities offer courses on DevOps, in which students learn about the tools and practices that enable rapid software development and deployment (Benni et al., 2019; Alves and Rocha, 2021; Fernandes et al., 2022; Verdicchio, 2023). Cloud infrastructure is a key part of DevOps, as it enables the automation of pipelines for building, testing, and deploying software. However, these courses mostly focus on the processes that make such a software lifecycle efficient, such as continuous integration and test orchestration, rather than the infrastructure that supports both these processes and the deployed applications.
@@ 3. Course Overview and Logistics
We designed a 10-week course entitled ”Cloud Infrastructure and Scalable Application Deployment” and offered it as an undergraduate elective at Stanford University in Winter 2024.
@@ 3.1. Learning Objectives
    Design cloud-native applications to take advantage of the elasticity, cost, and system administration benefits provided by cloud infrastructure.
@@
3.2. Class Composition
In total, 50 students were enrolled in the course, including 15 women and 7 students from underrepresented minority backgrounds. Most students were either juniors, seniors, or graduate students, although there were a limited number of underclassmen enrolled. Although the course was open to all students, most students were CS majors.
We intentionally required only programming maturity up to an introductory systems class and some experience using command-line tools in order to keep the course accessible to underclassmen. However, of the 34 students that completed an intro survey and remained enrolled in the class until the end, 76% had completed an operating systems course and 41% had completed a web or mobile development course. Additionally, 79% had some experience with backend or frontend web development, 68% had experience using an IaaS cloud provider, and 35% had experience with containerization. The first offering of the course was a pilot that more experienced students likely self-selected into, and we expect that future offerings will have a more diverse set of students.
@@ 3.3. Course Structure
Over 10 weeks, the course met for 80 minutes twice a week. Lectures covered the core areas of cloud application deployment, detailing the types of services offered by public cloud providers and the considerations behind selecting and deploying these services for a given application. We additionally discussed system and application design for scalability, security, and cost efficiency.
There were four hands-on programming assignments, each worth 15% of a student’s final grade, which we describe in Section 4.2. Assignments focused on systematically deploying a pre-built application to the cloud in modern best-practice patterns using IaC. We encouraged students to work in pairs on these assignments, and students were allowed to discuss conceptual strategies with other students in the class.
Lastly, the course featured a final project, worth 40% of students’ final grades, where students were given the opportunity to deploy an application of their own choosing to the cloud.
@@ 4. Course Content
Our course presented fundamental practices necessary for publishing and managing usable web applications on the public cloud as well as application design that takes advantage of cloud abstractions. Our lectures, assignments, and infrastructure for managing the class are open-source and publicly available online at https://infracourse.cloud.
@@ 4.1. Lecture Content
The first half of our course discussed the building blocks of cloud infrastructure that help in deploying web applications, including the basics of compute, network, and storage resources on the cloud, containers and container orchestration, and serverless and edge computing using Functions-as-a-Service (FaaS).
@@
Security was a significant focus area. We emphasized the shared responsibility model of cloud security, where providers take responsibility for securing the hardware and software underlying cloud platforms whereas developers take responsibility for how they use cloud services. We included a dedicated lecture on identity and access management (IAM) and its high attack surface in the context of modern serverless and containerized workloads running on the cloud. For all other lectures, where relevant, we included security considerations and best practices to minimize the attack surface of the services we discussed.
We covered all conceptual material and general concepts in lectures in a cloud provider-agnostic fashion, but demonstrated specific examples on Amazon Web Services (AWS) where required. We summarize the content covered in these lectures in Table 1.
@@ 4.1.1. Guest Lectures
We invited four guest lecturers from industry to provide insight and depth on select topics. All speakers were involved with leading the design and deployment of cloud-native applications at their respective employers. We summarize the content they covered in Table 1.
@@ 4 Continuous Integration & Deployment, Container Registries
Our course featured four hands-on technical assignments that were designed to give students experience with the cloud tools and services discussed in lectures. Our assignments focused on deploying web services to AWS, which we chose due to its predominant market share and the maturity of its cloud service offerings. AWS, via our institution’s account manager, additionally granted each student \$200 in credits for the duration of the course to cover resource costs. However, we emphasize that these assignments can be generalized to any cloud provider.
@@
Student completion rates were high for each assignment: 100% completed Assignment 1, 92% completed Assignments 2 and 3, and 90% completed Assignment 4.
@@ 4.3. Final Project
The final project tasked students with deploying a web application of their choosing to the public cloud. Students were allowed to choose either an application they designed themselves or an existing open-source project. The project was a way for students to cumulatively apply the concepts they learned without the scaffolding we provided in the assignments. In particular, it gave students the opportunity to architect a new deployment rather than merely instantiate cloud resources for an existing architecture.
Students were expected to tie together multiple cloud services using IaC to create an elastic deployment that could scale with demand and followed security best practices. Deliverables included the source code for the IaC used to deploy the application, a brief write-up of the deployment’s architecture and services used, and a link to the deployed application for evaluation.
@@ 5. Discussion
In total, 39 students filled out the survey. 72% of students stated they learned “a great deal” or “a lot” (a score of 4 or 5 on a 1-5 scale). Students broadly found value in the practical skills they learned: a few students reported that the course built up “actual confidence in my ability to use the cloud intelligently,” and that “this was the missing building block in building full-stack cloud awareness.” Students also overwhelmingly reported that they found the course useful for their future careers.
We also received feedback that the course was harder and more time-consuming than students expected. We believe this was partially due to the number of new concepts, services, and tools that students were exposed to, as well as general challenges common to any first offering of a course. However, many issues also stemmed from our miscalibration of assignment difficulty with the time we gave students to complete them, the scaffolding we provided in assignment starter code that prevented students from making more architectural choices, and state management issues with the AWS CDK that led to significant challenges in debugging deployments. We are currently in the process of addressing these issues based on student feedback for the next offering of the course.
@@ 5.1.1. Course infrastructure
To easily manage a course of 50 students with only three staff, we built a significant amount of custom infrastructure to support students given the variety of resources they interacted with throughout the course.
@@ 5.2.1. Building domain-specific debugging skills
During our pilot offering, we erred on the side of providing students with more scaffolding for IaC deployments. Unfortunately, this made it more difficult for students to practically apply application design concepts they learned in lecture. Students did interface with such concepts in instrumenting applications for observability and in some final projects that featured custom application logic. However, we recognize that we could have done more to let students explore these concepts in their assignments.
In the future, we intend to give students more architectural freedom for their submissions. For example, we can have students build further features for Yoctogram that take advantage of cloud-native constructs, such as image manipulation using functions-as-a-service. We can also envision grading some assignments based on performance and reliability rather than on the IaC itself, which may help in making assignments more open-ended without much additional overhead.
@@ 5.2.3. Adapting to students’ backgrounds
5.2.4. Account and billing logistics
Students worked on technical assignments using accounts they created themselves into which they could deposit the \$200 credit codes we provided. This required students to enter credit card information as a backup in case of overspend, but allowed students to keep unused credits after the term to use on personal projects.
The vast majority of students were able to avoid overspend as completing the assignments and final project in pairs allowed them to use a second AWS account in case funds in one account ran low. However, a few students did not properly clean up resources after the end of the term and thus saw charges on their credit cards. In these cases, we were able to work with AWS to waive the charges and get the resources shut down.
This process could be improved by having our institution provision student accounts using AWS Organizations and IAM Identity Center, which transfers overspend liability from the student to the institution. Managing such infrastructure would increase the course staff’s workload, but it prevents students from incurring charges and would also allow staff to more easily troubleshoot student account issues. We also plan to cover basic billing topics earlier in the course and to ensure students set billing limits on their accounts to prevent overspend.
@@ 6. Conclusion
In this work, we presented our experience teaching an undergraduate course on cloud infrastructure and scalable application deployment. Students learned practical skills related to deploying and managing cloud infrastructure, as well as application design that takes advantage of managed cloud services. Our course featured a number of hands-on assignments that gave students experience with modern, best-practice concepts such as IaC, containerization, and serverless computing. While there are several improvements we can make to future offerings, both students and course staff felt that the course was successful in building student fluency with cloud fundamentals and tools useful for launching scalable applications and for further career growth. We hope that our experiences and public course materials can be useful to educators interested in addressing the gap in practical cloud computing education for undergraduate students.